A major airline company is concerned that its proportion of late arrivals has substantially increased in the past month. Historical data shows that on the average 18% of the company airplanes have arrived late. In a random sample of 1,240 airplanes, 310 airplanes have arrived late. If we are conducting a hypothesis test of a single proportion to determine if the proportion of late arrivals has increased: What is the value of the calculated test statistic?
Answer #1

Sample proportion = 310 / 1240 = 0.25

Test statistics

z = - p / sqrt( p (1 - p ) / n)

= 0.25 - 0.18 / sqrt( 0.18 * 0.82 / 1240)

= 6.42